OpenStudy (anonymous):

One pair of corresponding sides of two similar polygons measures 12 and 15. The perimeter of the smaller polygon is 30. Find the perimeter of the larger.

pooja195 (pooja195):

One pair of corresponding sides of two similar polygons measures 12 and 15. The perimeter of the smaller polygon is 30. Find the perimeter of the larger. Again, the ratio is 15:12 = 5/4. (Notice I'm picking which number to go on top of the fraction by checking whether I want to calculate a measurement on the larger or the smaller polygon.) 30 * 5/4 = 150/4 = 37.5